OpenStudy (hang254):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 25. What is the length of the altitude? 41 50 25 20

OpenStudy (hang254):

@Calcmathlete

OpenStudy (anonymous):

Solve this proportion: \[\frac{16}{x} = \frac{x}{25}\]

OpenStudy (hang254):

400=x

OpenStudy (hang254):

oh, i get it

OpenStudy (hang254):

x=25. thanks again

OpenStudy (anonymous):

Not quite. x = 20 x^2 = 400 x = √400 = 20

OpenStudy (hang254):

whaaaaaat. but 16/25=.8 and 20/25=.8

OpenStudy (hang254):

oh, i just said the wrong thing. my bad.